54. A bicyclist rode into the county for 5 h. In returning, her speed was 5 mi/h faster and the trip took 4h. What was her speed each way.
Answer by rajagopalan(174)

Let onward speed = x
Let distance = d
on ward time = d/x = 5 (given) ..equation 1
..
return
speed = x+5
distance = d
return journey time = d/(x+5)= 4 (given)..equation 2
..
from eq 1 .. we get d/x=5, 5x=d
substitute d=5x in eqn 2
d/(x+5) = 4
5x/(x+5)=4
5x = 4(x+5)
5x = 4x+20
5x-4x = 20
x = 20.
Onward speed = 20
Return speed = 25 ..
good luck..